OpenStudy (anonymous):

This is a simple one but I don't have a calculator and none of the online ones I see are helping much.. 4/-3 (that's four divided by negative three) *7. (Multiplied by seven). Could someone just shoot me the answer?

OpenStudy (zehanz):

\[\frac{ 4 }{ -3 }\cdot 7=-\frac{ 4 }{ 3 }\cdot \frac{ 7 }{ 1 }=-\frac{ 4 \cdot 7 }{ 3 \cdot 1 }=-\frac{ 28 }{ 3 }\]If you want, you can write it as -9.333333333...

OpenStudy (anonymous):

Thank you :) this helped a lot. :D

OpenStudy (zehanz):

YW!